An arrow weighing 7 g is mounted on a square piece of wood weighing 1 kg
penetrated to a depth of 8 cm. Then place the square stick with the arrow on a non-abrasive horizontal surface
How many centimeters deep will the arrow penetrate when it is placed and fired again?
1
Expert's answer
2020-05-04T12:46:45-0400
"\Delta E\backsim l"

"\Delta E_1=0.5v^2\frac{mM}{m+M}, \Delta E_2=0.5mv^2"

"l_2=l_1\frac{m+M}{M}=8\frac{0.007+1}{1}=8.056\ cm"